OpenStudy (anonymous):

Solve the equation. sqrt 2z+3=1- sqrt z+1= z=

OpenStudy (anonymous):

square both sides...

OpenStudy (anonymous):

sqrt 2z+3=1- sqrt z+1= sqrt( 2z+3)=1- sqrt (z+1)...squaring both sides to get 2z+3=1-2sqrt(z+1)+z+1 2z-z+3-2=2sqrt(z+1) z+1=2sqrt(z+1),,squaring both sides again to get z^2 +2z+1=4(z+1) z^2 +2z+1=4z+4 z^2-2z-3=0,factoring (z-3)(z+1)=0 z=3 and z=-1 ans